OpenStudy (anonymous):

Which point lies on a circle with a radius of 5 units and center at P(6, 1)?

OpenStudy (anonymous):

Q(1, 11) R(2, 4) S(4, -4) T(9, -2)

OpenStudy (anonymous):

possible answers

OpenStudy (mosaic):

The general equation of a circle is: (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center and r is the radius. Find the equation of the circle first and try each answer to find the equation that satisfies.

OpenStudy (anonymous):

im completely lost lol

OpenStudy (mosaic):

The center is (6,1). Therefore, h = 6, y = 1. The radius r is given as 5. So the equation of the circle is: (x - 6)^2 + (y - 1)^2 = 5^2 = 25 To see if the first choice, Q(1, 11), is on the circle, substitute x = 1, y = 11 in (x - 6)^2 + (y - 1)^2 = 25 and see if it satisfies the equation. If not, try the next choice.

OpenStudy (mosaic):

h = 6, k = 1